# {textblock*}

### v3

@@ 5.2.2 Performance on Enzymes, Proteins and DD against state of the art graph neural networks
MLP 4 55.833$\pm$3.516 93.062$\pm$7.551 332.30
vanilla GCN 4 65.833$\pm$4.610 97.688$\pm$3.064 343.00
GraphSage 4 65.000$\pm$4.944 100.000$\pm$0.000 294.20
MoNet 4 63.000$\pm$8.090 95.229$\pm$5.864 333.70
GAT 4 68.500$\pm$5.241 100.000$\pm$0.000 299.30
GatedGCN 4 65.667$\pm$4.899 99.979$\pm$0.062 316.80
GIN 4 65.333$\pm$6.823 100.000$\pm$0.000 402.10
RingGNN 2 18.667$\pm$1.795 20.104$\pm$2.166 337.30
3WLGNN 3 61.000$\pm$6.799 98.875$\pm$1.571 381.80
HGNN* - 51.300$\pm$6.100 - -
H2H-GCN* - 61.300$\pm$4.900 - -
LGIN Variable $c$(Ours) 3 67.380$\pm$4.100 100.000$\pm$0.000 500.00
LGIN Fixed $c$(Ours) 3 71.500$\pm$1.600 100.000$\pm$0.000 500.00
@@
MLP 4 72.239$\pm$3.854 73.816$\pm$1.015 371.80
vanilla GCN 4 72.758$\pm$4.083 100.000$\pm$0.000 266.70
GraphSage 4 73.433$\pm$3.429 75.289$\pm$2.419 267.20
MoNet 4 71.736$\pm$3.365 81.003$\pm$2.593 252.60
GAT 4 75.900$\pm$3.824 95.851$\pm$2.575 201.60
GatedGCN 4 72.918$\pm$2.090 82.796$\pm$2.242 300.70
GIN 4 71.910$\pm$3.873 99.851$\pm$0.136 275.70
RingGNN 2 - - -
3WLGNN 3 - - -
HGNN* - 75.800$\pm$3.300 - -
H2H-GCN* - 78.200$\pm$3.300 - -
LGIN Variable $c$(Ours) 3 73.730$\pm$3.200 100.000$\pm$0.000 297.00
LGIN Fixed $c$(Ours) 3 70.300$\pm$3.500 100.000$\pm$0.000 281.00
@@
MLP 4 75.644$\pm$2.681 79.847$\pm$1.551 244.20
vanilla GCN 4 76.098$\pm$2.406 81.387$\pm$2.451 350.90
GraphSage 4 75.289$\pm$2.419 85.182$\pm$3.489 245.40
MoNet 4 76.452$\pm$2.898 78.206$\pm$0.548 306.10
GAT 4 76.277$\pm$2.410 83.186$\pm$2.000 344.10
GatedGCN 4 76.363$\pm$2.904 79.471$\pm$0.695 293.80
GIN 4 74.117$\pm$3.357 75.351$\pm$1.267 420.90
RingGNN 2 67.564$\pm$7.551 67.564$\pm$7.551 450.10
3WLGNN 3 61.712$\pm$4.859 62.427$\pm$4.548 211.40
HGNN* - 73.700$\pm$2.300 - -
H2H-GCN* - 74.400$\pm$3.000 - -
LGIN Variable $c$(Ours) 3 79.100$\pm$4.400 81.750$\pm$10.030 1000.00
LGIN Fixed $c$(Ours) 3 76.900$\pm$2.500 84.180$\pm$5.530 700.00
Table 2 presents the detailed performance analysis across the Proteins, Enzymes, and DD datasets. We evaluate multiple models, including Euclidean GNNs and hyperbolic GNN baselines(HGNN and H2H-GCN), alongside our proposed methods LGIN Variable and LGIN Fixed. On the Enzymes dataset, our LGIN Fixed model achieves the highest test accuracy of 71.500%, significantly outperforming the hyperbolic baselines HGNN(51.3%) and H2H-GCN(61.3%), and other Euclidean models. GAT achieves the second best performance (68.500%). The LGIN Variable variant follows closely with 67.380%. For the DD dataset, the hyperbolic baseline H2H-GCN achieves the highest test accuracy of 78.2%. GAT is the second best at 75.900%. HGNN performs very close to GAT (75.8%). Our LGIN Variable model achieves a test accuracy of 73.730%, performing comparably to HGNN and better than many Euclidean baselines, and is the fourth best overall on this dataset. LGIN Fixed has lower performance on DD (70.300%). In the Proteins dataset, our LGIN Variable model attains the highest accuracy at 79.100%, surpassing both LGIN Fixed (76.900%) and the hyperbolic baselines HGNN(73.7%) and H2H-GCN(74.4%). H2H-GCN achieves the third best performance on this dataset. Overall, the results demonstrate that LGIN models remain highly competitive across these datasets. LGIN Fixed is the top performer on Enzymes, while LGIN Variable leads on Proteins. Although H2H-GCN is the best on DD, LGIN Variable and Fixed show strong performance on Proteins and Enzymes respectively, highlighting LGIN’s ability to model complex graph structures effectively within the Lorentzian geometry, competing well against both Euclidean and other hyperbolic methods.
5.2.3 Performance on MoleculeNet Datasets
Table 3: Analysis of our framework’s performance on MoleculeNet baselines. The best score is marked in bold and the second best score is underlined. GIN models have been marked with an asterisk to indicate that our proposed model significantly outperforms GIN baselines. Hyphen indicates OOM on an 8GB M2 Macbook Pro.
@@ 5.2.3 Performance on MoleculeNet Datasets
Figure 1: Abalation Study on Curvature Type and Parallel Transport on BBBP
Table 3 presents the performance analysis of our framework on the MoleculeNet datasets, comparing it against various baseline models, including GAT, GIN, and DGN, as well as their enhanced versions incorporating ECFP and MFP features. Results are reported across three datasets: BACE, BBBP, and HIV.
Our method achieves state-of-the-art performance on multiple datasets. Specifically, on the BBBP dataset, LGIN Fixed $c$ achieves the highest score of 91.54, outperforming all other models, while LGIN Variable $c$ secures the second-best score of 89.08. On the BACE dataset, DGCL achieves the best performance with 91.48, closely followed by LGIN Variable $c$ with 90.30, surpassing the performance of GIN-based models such as GIN-ECFP (82.84*) and GIN-MFP (82.03*). Similarly, for the HIV dataset, DGCL attains the highest score of 81.49, with GAT-MFP achieving the second-best score of 76.47.
Furthermore, our LGIN models consistently outperform the baseline GIN models across all datasets, demonstrating the effectiveness of leveraging Lorentzian geometry for molecular representation learning. These results highlight the robustness of our approach in capturing complex molecular structures, thereby achieving superior performance in molecular property prediction tasks.
5.2.4 Empirical Analysis on Curvature and Parallel Transport
$\mathbf{x}^{\mathcal{L}^{k}}=\text{LT}^{k}(\exp_{o}^{c}((1+\epsilon^{k}).\log_{o}^{c}(\mathbf{x}^{\mathcal{L}^{k-1}})+\log_{o}^{c}(T(\mathbf{x}^{\mathcal{L}^{k-1}}))))$ (21)
figure 1 represents the AUC scores for curvature and parallel transport modes while provides a direct comparison between the results of the two update rules. As noted, cases where parallel transport is employed to the first term of eq.21 perform considerably better. We hypothesize that this is probably due to the fact that parallel transport aligns embeddings which were otherwise distorted due to aggregation, despite all transformations taking place with respect to the origin. With respect to curvature type, there is no significant difference between fixed and variable curvature models, in line with evaluations described in previous sections and (Zhang et al., 2021a).
6 Conclusion and Limitations
In this paper, we address the need for simple yet effective graph neural networks on hyperbolic manifolds capable of capturing both semantic and structural features. We propose the Lorentzian Graph Isomorphic Network (LGIN), which, to the best of our knowledge, is the first framework to extend graph isomorphism tests to hyperbolic spaces with variable and fixed curvature. Through extensive experiments, we demonstrate that LGIN consistently matches or surpasses state-of-the-art models across multiple graph benchmarks. Additionally, we provide a detailed explanation of our update rule, including rigorous derivations for each step, thereby establishing a framework for an approximate solution for the expressivity bottleneck on hyperbolic spaces.
However, our analysis does not offer a comprehensive framework for hyperbolic graph neural networks that directly map between hyperbolic spaces without relying on tangent space projections. Further investigation is required to develop such transformations and fully explore the potential of these models, which we leave as a direction for future research.
@@ 6 Conclusion and Limitations
Let the transformation be denoted by $F(x^{\mathcal{L}})=\mathbf{W}\otimes_{c}^{\mathcal{L}}x^{\mathcal{L}}$. We can decompose $F$ into a sequence of maps: $F=f_{5}\circ f_{4}\circ f_{3}\circ f_{2}\circ f_{1}$, where:
  1. 1.
    $f_{1}:\mathcal{H}^{n}\to T_{o}\mathcal{H}^{n}$, $f_{1}(x^{\mathcal{L}})=\log_{o}^{c}(x^{\mathcal{L}})$. The logarithmic map centered at the origin $o$ of the hyperboloid model $\mathcal{H}^{n}$ is a global diffeomorphism from $\mathcal{H}^{n}$ (the upper sheet) to $T_{o}\mathcal{H}^{n}$. As a diffeomorphism, $f_{1}$ is bijective and thus injective.
  2. 2.
    $f_{2}:T_{o}\mathcal{H}^{n}\to\mathbb{R}^{n}$, $f_{2}({\bm{c}}{v})={\bm{c}}{v}_{[1:n]}$. In the hyperboloid model embedded in $\mathbb{R}^{n+1}$, the tangent space at the origin $o=(1,0,\dots,0)$ consists of vectors ${\bm{c}}{v}=(v_{0},v_{1},\dots,v_{n})\in\mathbb{R}^{n+1}$ such that $\langle o,{\bm{c}}{v}\rangle_{\eta}=-1\cdot v_{0}+\sum_{i=1}^{n}0\cdot v_{i}=-v_{0}=0$. Thus $T_{o}\mathcal{H}^{n}=\{(0,v_{1},\dots,v_{n})\mid v_{i}\in\mathbb{R}\}$. The map $f_{2}$ extracts the spatial components $(v_{1},\dots,v_{n})$, providing an isomorphism between $T_{o}\mathcal{H}^{n}$ and $\mathbb{R}^{n}$. As an isomorphism, $f_{2}$ is bijective and thus injective.
  3. 3.
    $f_{3}:\mathbb{R}^{n}\to\mathbb{R}^{d}$, $f_{3}({\bm{c}}{y})=\mathbf{W}{\bm{c}}{y}$. This is a standard linear transformation. A linear transformation $f_{3}$ is injective if and only if its null space, $\ker(\mathbf{W})=\{{\bm{c}}{y}\in\mathbb{R}^{n}\mid\mathbf{W}{\bm{c}}{y}={\bm{c}}{0}\}$, contains only the zero vector. By the Rank-Nullity Theorem ($n=\dim(\ker(\mathbf{W}))+\text{rank}(\mathbf{W})$), this is equivalent to $\text{rank}(\mathbf{W})=n$. The maximum possible rank for $\mathbf{W}\in\mathbb{R}^{d\times n}$ is $\min(d,n)$. Thus, having rank $n$ requires $d\geq n$. $f_{3}$ is injective if and only if $\mathbf{W}$ has full column rank ($n$).
  4. 4.
    $f_{4}:\mathbb{R}^{d}\to T_{o}\mathcal{H}^{d}$, $f_{4}({\bm{c}}{z})=(0,{\bm{c}}{z})$. Similar to $f_{2}$, this map takes a vector ${\bm{c}}{z}\in\mathbb{R}^{d}$ and forms a vector $(0,z_{1},\dots,z_{d})\in\mathbb{R}^{d+1}$, which lies in $T_{o}\mathcal{H}^{d}$. This map is an isomorphism between $\mathbb{R}^{d}$ and $T_{o}\mathcal{H}^{d}$. As an isomorphism, $f_{4}$ is bijective and thus injective.
  5. 5.
    $f_{5}:T_{o}\mathcal{H}^{d}\to\mathcal{H}^{d}$, $f_{5}({\bm{c}}{u})=\exp_{o}^{c}({\bm{c}}{u})$. The exponential map centered at the origin $o$ is a global diffeomorphism from $T_{o}\mathcal{H}^{d}$ to $\mathcal{H}^{d}$ (the upper sheet). As a diffeomorphism, $f_{5}$ is bijective and thus injective.
@@
To demonstrate this explicitly, assume $F(x_{1}^{\mathcal{L}})=F(x_{2}^{\mathcal{L}})$ for $x_{1}^{\mathcal{L}},x_{2}^{\mathcal{L}}\in\mathcal{H}^{n}$.
$\exp_{o}^{c}\left(0,\mathbf{W}\left(\log_{o}^{c}x_{1}^{\mathcal{L}}\right)_{[1:n]}\right)=\exp_{o}^{c}\left(0,\mathbf{W}\left(\log_{o}^{c}x_{2}^{\mathcal{L}}\right)_{[1:n]}\right)$
Since $f_{5}=\exp_{o}^{c}$ is injective, we have:
$\left(0,\mathbf{W}\left(\log_{o}^{c}x_{1}^{\mathcal{L}}\right)_{[1:n]}\right)=\left(0,\mathbf{W}\left(\log_{o}^{c}x_{2}^{\mathcal{L}}\right)_{[1:n]}\right)$
@@
$\mathbf{W}\left(\log_{o}^{c}x_{1}^{\mathcal{L}}\right)_{[1:n]}=\mathbf{W}\left(\log_{o}^{c}x_{2}^{\mathcal{L}}\right)_{[1:n]}$
@@
Let ${\bm{c}}{y}_{1}=\left(\log_{o}^{c}x_{1}^{\mathcal{L}}\right)_{[1:n]}$ and ${\bm{c}}{y}_{2}=\left(\log_{o}^{c}x_{2}^{\mathcal{L}}\right)_{[1:n]}$. Both ${\bm{c}}{y}_{1},{\bm{c}}{y}_{2}\in\mathbb{R}^{n}$. The equation is $\mathbf{W}{\bm{c}}{y}_{1}=\mathbf{W}{\bm{c}}{y}_{2}$. If $\mathbf{W}$ has full column rank, the linear map $f_{3}({\bm{c}}{y})=\mathbf{W}{\bm{c}}{y}$ is injective, which implies ${\bm{c}}{y}_{1}={\bm{c}}{y}_{2}$.
$\left(\log_{o}^{c}x_{1}^{\mathcal{L}}\right)_{[1:n]}=\left(\log_{o}^{c}x_{2}^{\mathcal{L}}\right)_{[1:n]}$
@@
Since $f_{2}^{-1}$ maps $\mathbb{R}^{n}$ back to $T_{o}\mathcal{H}^{n}$ injectively, this implies the equality of the tangent vectors:
$\log_{o}^{c}x_{1}^{\mathcal{L}}=\log_{o}^{c}x_{2}^{\mathcal{L}}$
Finally, since $f_{1}=\log_{o}^{c}$ is injective, this implies:
$x_{1}^{\mathcal{L}}=x_{2}^{\mathcal{L}}$
@@
Thus, if $\mathbf{W}$ has full column rank, $F(x_{1}^{\mathcal{L}})=F(x_{2}^{\mathcal{L}})$ implies $x_{1}^{\mathcal{L}}=x_{2}^{\mathcal{L}}$, proving $F$ is injective.
@@
Conversely, assume $\mathbf{W}$ does not have full column rank. Since $\mathbf{W}\in\mathbb{R}^{d\times n}$, this means $\text{rank}(\mathbf{W})<n$. By the Rank-Nullity Theorem, $\dim(\ker(\mathbf{W}))=n-\text{rank}(\mathbf{W})>0$. This means there exists a non-zero vector ${\bm{c}}{v}\in\mathbb{R}^{n}$ such that $\mathbf{W}{\bm{c}}{v}={\bm{c}}{0}$. Let ${\bm{c}}{u}=f_{2}^{-1}({\bm{c}}{v})\in T_{o}\mathcal{H}^{n}$. Since $f_{2}^{-1}$ is injective and ${\bm{c}}{v}\neq{\bm{c}}{0}$, ${\bm{c}}{u}\neq{\bm{c}}{0}$. Let $x^{\mathcal{L}}=f_{1}^{-1}({\bm{c}}{u})\in\mathcal{H}^{n}$. Since $f_{1}^{-1}$ is injective and ${\bm{c}}{u}\neq{\bm{c}}{0}$, $x^{\mathcal{L}}\neq o$. Consider $F(x^{\mathcal{L}})$ and $F(o)$. $F(x^{\mathcal{L}})=f_{5}(f_{4}(f_{3}(f_{2}(f_{1}(x^{\mathcal{L}})))))=f_{5}(f_{4}(f_{3}({\bm{c}}{u})))=f_{5}(f_{4}(\mathbf{W}{\bm{c}}{v}))=f_{5}(f_{4}({\bm{c}}{0}))$. $f_{4}({\bm{c}}{0})=(0,{\bm{c}}{0})\in T_{o}\mathcal{H}^{d}$. $F(x^{\mathcal{L}})=\exp_{o}^{c}(0,{\bm{c}}{0})$. The exponential map of the zero vector in the tangent space at $p$ is always $p$. So, $\exp_{o}^{c}(0,{\bm{c}}{0})=o\in\mathcal{H}^{d}$. Now consider $F(o)$: $F(o)=\exp_{o}^{c}\left(0,\mathbf{W}\left(\log_{o}^{c}o\right)_{[1:n]}\right)$. $\log_{o}^{c}o$ is the zero vector in $T_{o}\mathcal{H}^{n}$, which is $(0,{\bm{c}}{0})\in\mathbb{R}^{n+1}$. $\left(\log_{o}^{c}o\right)_{[1:n]}={\bm{c}}{0}\in\mathbb{R}^{n}$. $F(o)=\exp_{o}^{c}(0,\mathbf{W}{\bm{c}}{0})=\exp_{o}^{c}(0,{\bm{c}}{0})=o\in\mathcal{H}^{d}$. So $F(x^{\mathcal{L}})=o$ and $F(o)=o$. Since $x^{\mathcal{L}}\neq o$, we have found two distinct points $x^{\mathcal{L}},o\in\mathcal{H}^{n}$ that map to the same point $o\in\mathcal{H}^{d}$. Thus, $F$ is not injective if $\mathbf{W}$ does not have full column rank.

### v4

@@ 4.1.1. Evaluations against Graph Isomorphic Network
MLP 4 55.833$\pm$3.516 93.062$\pm$7.551
vanilla GCN 4 65.833$\pm$4.610 97.688$\pm$3.064
GraphSage 4 65.000$\pm$4.944 100.000$\pm$0.000
MoNet 4 63.000$\pm$8.090 95.229$\pm$5.864
GAT 4 68.500$\pm$5.241 100.000$\pm$0.000
GatedGCN 4 65.667$\pm$4.899 99.979$\pm$0.062
GIN 4 65.333$\pm$6.823 100.000$\pm$0.000
RingGNN 2 18.667$\pm$1.795 20.104$\pm$2.166
3WLGNN 3 61.000$\pm$6.799 98.875$\pm$1.571
HGNN* - 51.300$\pm$6.100 -
H2H-GCN* - 61.300$\pm$4.900 -
LGIN Variable $c$(Ours) 3 67.380$\pm$4.100 100.000$\pm$0.000
LGIN Fixed $c$(Ours) 3 71.500$\pm$1.600 100.000$\pm$0.000
@@
MLP 4 72.239$\pm$3.854 73.816$\pm$1.015
vanilla GCN 4 72.758$\pm$4.083 100.000$\pm$0.000
GraphSage 4 73.433$\pm$3.429 75.289$\pm$2.419
MoNet 4 71.736$\pm$3.365 81.003$\pm$2.593
GAT 4 75.900$\pm$3.824 95.851$\pm$2.575
GatedGCN 4 72.918$\pm$2.090 82.796$\pm$2.242
GIN 4 71.910$\pm$3.873 99.851$\pm$0.136
RingGNN 2 - -
3WLGNN 3 - -
HGNN* - 75.800$\pm$3.300 -
H2H-GCN* - 78.200$\pm$3.300 -
LGIN Variable $c$(Ours) 3 73.730$\pm$3.200 100.000$\pm$0.000
LGIN Fixed $c$(Ours) 3 70.300$\pm$3.500 100.000$\pm$0.000
@@
MLP 4 75.644$\pm$2.681 79.847$\pm$1.551
vanilla GCN 4 76.098$\pm$2.406 81.387$\pm$2.451
GraphSage 4 75.289$\pm$2.419 85.182$\pm$3.489
MoNet 4 76.452$\pm$2.898 78.206$\pm$0.548
GAT 4 76.277$\pm$2.410 83.186$\pm$2.000
GatedGCN 4 76.363$\pm$2.904 79.471$\pm$0.695
GIN 4 74.117$\pm$3.357 75.351$\pm$1.267
RingGNN 2 67.564$\pm$7.551 67.564$\pm$7.551
3WLGNN 3 61.712$\pm$4.859 62.427$\pm$4.548
HGNN* - 73.700$\pm$2.300 -
H2H-GCN* - 74.400$\pm$3.000 -
LGIN Variable $c$(Ours) 3 79.100$\pm$4.400 81.750$\pm$10.030
LGIN Fixed $c$(Ours) 3 76.900$\pm$2.500 84.180$\pm$5.530
Table 1 presents a performance comparison of LGIN variants against powerful Euclidean baselines, including GIN variants, WL kernels, and others, across MUTAG, PROTEINS, PTC, and NCI1 datasets. The proposed LGIN variants demonstrate competitive and often superior performance. Specifically, LGIN with Fixed Curvature achieves the highest accuracy on the MUTAG (90.6 $\pm$ 3.9) and PTC (68.4 $\pm$ 5.2) datasets. LGIN with Variable Curvature attains the best performance on the PROTEINS dataset (79.1 $\pm$ 4.4), showing statistically significant improvement ($p<0.1$). While the WL subtree kernel leads on NCI1 (86.0 $\pm$ 1.8) and PATCHYSAN is competitive on MUTAG (92.6 $\pm$ 4.2), LGIN variants consistently rank among the top performers. These results underscore LGIN’s ability to effectively leverage Lorentzian geometry and incorporated curvature for enhanced representation learning and discrimination across graphs with varying structures, demonstrating the benefits of adapting isomorphic principles to non-Euclidean spaces.
4.1.2. Performance on Enzymes, Proteins and DD against state of the art graph neural networks
Table 2 details performance across ENZYMES, DD, and PROTEINS, evaluating LGIN variants, Euclidean GNNs, and other HGNNs. Performance on ENZYMES, DD, and PROTEINS is shown in Table 2. On ENZYMES, our LGIN Fixed model achieves the highest accuracy (71.5%), significantly outperforming hyperbolic baselines, with GAT second. For DD, H2H-GCN leads (78.2%), followed by GAT and HGNN; our LGIN Variable (73.70%) is competitive. In PROTEINS, our LGIN Variable attains the highest accuracy (79.1%), surpassing LGIN Fixed and other hyperbolic baselines. Overall, LGIN variants are highly competitive across these datasets. LGIN Fixed is best on ENZYMES, LGIN Variable on PROTEINS. While H2H-GCN is best on DD, LGIN variants perform strongly, highlighting their effectiveness on complex graph structures.
4.1.3. Performance on MoleculeNet Datasets
Table 3. Analysis of our framework’s performance on MoleculeNet baselines. The best score is marked in bold and the second best score is underlined. GIN models have been marked with an asterisk to indicate that our proposed model significantly outperforms GIN baselines. Hyphen indicates OOM on an 8GB M2 Macbook Pro.
@@ 4.1.3. Performance on MoleculeNet Datasets
Table 3 presents performance on the BACE, BBBP, and HIV MoleculeNet datasets, comparing LGIN variants against GAT, GIN, DGN, DGCL, and enhanced versions with ECFP/MFP features. Our method achieves competitive and state-of-the-art results. Specifically, LGIN Fixed $c$ attains the highest score on the BBBP dataset (91.54), with LGIN Variable $c$ second (89.08). On BACE, DGCL leads (91.48), closely followed by LGIN Variable $c$ (90.30), surpassing GIN-based models (GIN-ECFP 82.84*, GIN-MFP 82.03*). For HIV, DGCL is best (81.49), with GAT-MFP second (76.47). LGIN variants consistently outperform baseline GIN models across all datasets. These results demonstrate LGIN’s effectiveness in leveraging Lorentzian geometry for molecular representation learning and achieving superior performance in molecular property prediction.
5. Conclusion and Limitations
In this paper, we address the need for simple yet effective graph neural networks on hyperbolic manifolds capable of capturing both semantic and structural features. We propose the Lorentzian Graph Isomorphic Network (LGIN), which, to the best of our knowledge, is the first framework to extend graph isomorphism tests to hyperbolic spaces with variable and fixed curvature. However, our analysis does not offer a comprehensive framework for hyperbolic graph neural networks that directly map between hyperbolic spaces without relying on tangent space projections. Further investigation is required to develop such transformations and fully explore the potential of these models, which we leave as a direction for future research.
@@ 5. Conclusion and Limitations
Let the transformation be denoted by $F(x)=\mathcal{W}(x)$. We can decompose $F$ into a sequence of maps: $F=f_{5}\circ f_{4}\circ f_{3}\circ f_{2}\circ f_{1}$, where:
  1. (1)
    $f_{1}:H^{n}\to T_{o_{n}}H^{n}$, $f_{1}(x)=\log_{o_{n}}^{c}(x)$. The logarithmic map centered at the origin $o_{n}$ of the hyperboloid model $H^{n}$ is a global diffeomorphism from $H^{n}$ (the upper sheet) to $T_{o_{n}}H^{n}$. As a diffeomorphism, $f_{1}$ is bijective and thus injective.
  2. (2)
    $f_{2}:T_{o_{n}}H^{n}\to\mathbb{R}^{n}$, $f_{2}(\mathbf{v})=\mathbf{v}_{[1:n]}$. In the hyperboloid model embedded in $\mathbb{R}^{n+1}$, the origin is $o_{n}=(1,0,\dots,0)$. The tangent space at the origin $o_{n}$ consists of vectors $\mathbf{v}=(v_{0},v_{1},\dots,v_{n})\in\mathbb{R}^{n+1}$ such that $\langle o_{n},\mathbf{v}\rangle_{\eta}=-1\cdot v_{0}+\sum_{i=1}^{n}0\cdot v_{i}=-v_{0}=0$. Thus $T_{o_{n}}H^{n}=\{(0,v_{1},\dots,v_{n})\mid v_{i}\in\mathbb{R}\}$. The map $f_{2}$ extracts the spatial components $(v_{1},\dots,v_{n})$, providing an isomorphism between $T_{o_{n}}H^{n}$ and $\mathbb{R}^{n}$. As an isomorphism, $f_{2}$ is bijective and thus injective.
  3. (3)
    $f_{3}:\mathbb{R}^{n}\to\mathbb{R}^{d}$, $f_{3}(\mathbf{y})=\mathbf{W}\mathbf{y}$. This is a standard linear transformation. A linear transformation $f_{3}$ is injective if and only if its null space, $\ker(\mathbf{W})=\{\mathbf{y}\in\mathbb{R}^{n}\mid\mathbf{W}\mathbf{y}=\mathbf{0}\}$, contains only the zero vector. By the Rank-Nullity Theorem ($n=\dim(\ker(\mathbf{W}))+\text{rank}(\mathbf{W})$), this is equivalent to $\text{rank}(\mathbf{W})=n$. The maximum possible rank for $\mathbf{W}\in\mathbb{R}^{d\times n}$ is $\min(d,n)$. Thus, having rank $n$ requires $d\geq n$. $f_{3}$ is injective if and only if $\mathbf{W}$ has full column rank ($n$).
  4. (4)
    $f_{4}:\mathbb{R}^{d}\to T_{o_{d}}H^{d}$, $f_{4}(\mathbf{z})=(0,\mathbf{z})$. Similar to $f_{2}$, this map takes a vector $\mathbf{z}\in\mathbb{R}^{d}$ and forms a vector $(0,z_{1},\dots,z_{d})\in\mathbb{R}^{d+1}$, which lies in $T_{o_{d}}H^{d}$. This map is an isomorphism between $\mathbb{R}^{d}$ and $T_{o_{d}}H^{d}$. As an isomorphism, $f_{4}$ is bijective and thus injective.
  5. (5)
    $f_{5}:T_{o_{d}}H^{d}\to H^{d}$, $f_{5}(\mathbf{u})=\exp_{o_{d}}^{c}(\mathbf{u})$. The exponential map centered at the origin $o_{d}$ is a global diffeomorphism from $T_{o_{d}}H^{d}$ to $H^{d}$ (the upper sheet). As a diffeomorphism, $f_{5}$ is bijective and thus injective.
@@
To demonstrate this explicitly, assume $F(x_{1})=F(x_{2})$ for $x_{1},x_{2}\in H^{n}$.
$\exp_{o_{d}}^{c}\left(0,\mathbf{W}\left(\log_{o_{n}}^{c}x_{1}\right)_{[1:n]}\right)=\exp_{o_{d}}^{c}\left(0,\mathbf{W}\left(\log_{o_{n}}^{c}x_{2}\right)_{[1:n]}\right)$
Since $f_{5}=\exp_{o_{d}}^{c}$ is injective, we have:
$\left(0,\mathbf{W}\left(\log_{o_{n}}^{c}x_{1}\right)_{[1:n]}\right)=\left(0,\mathbf{W}\left(\log_{o_{n}}^{c}x_{2}\right)_{[1:n]}\right)$
@@
$\mathbf{W}\left(\log_{o_{n}}^{c}x_{1}\right)_{[1:n]}=\mathbf{W}\left(\log_{o_{n}}^{c}x_{2}\right)_{[1:n]}$
@@
Let $\mathbf{y}_{1}=\left(\log_{o_{n}}^{c}x_{1}\right)_{[1:n]}$ and $\mathbf{y}_{2}=\left(\log_{o_{n}}^{c}x_{2}\right)_{[1:n]}$. Both $\mathbf{y}_{1},\mathbf{y}_{2}\in\mathbb{R}^{n}$. The equation is $\mathbf{W}\mathbf{y}_{1}=\mathbf{W}\mathbf{y}_{2}$. If $\mathbf{W}$ has full column rank, the linear map $f_{3}(\mathbf{y})=\mathbf{W}\mathbf{y}$ is injective, which implies $\mathbf{y}_{1}=\mathbf{y}_{2}$.
$\left(\log_{o_{n}}^{c}x_{1}\right)_{[1:n]}=\left(\log_{o_{n}}^{c}x_{2}\right)_{[1:n]}$
@@
Since $f_{2}^{-1}$ maps $\mathbb{R}^{n}$ back to $T_{o_{n}}H^{n}$ injectively, this implies the equality of the tangent vectors:
$\log_{o_{n}}^{c}x_{1}=\log_{o_{n}}^{c}x_{2}$
Finally, since $f_{1}=\log_{o_{n}}^{c}$ is injective, this implies:
$x_{1}=x_{2}$
@@
Thus, if $\mathbf{W}$ has full column rank, $F(x_{1})=F(x_{2})$ implies $x_{1}=x_{2}$, proving $F$ is injective.
@@
Conversely, assume $\mathbf{W}$ does not have full column rank. Since $\mathbf{W}\in\mathbb{R}^{d\times n}$, this means $\text{rank}(\mathbf{W})<n$. By the Rank-Nullity Theorem, $\dim(\ker(\mathbf{W}))=n-\text{rank}(\mathbf{W})>0$. This means there exists a non-zero vector $\mathbf{v}\in\mathbb{R}^{n}$ such that $\mathbf{W}\mathbf{v}=\mathbf{0}$. Let $\mathbf{u}=f_{2}^{-1}(\mathbf{v})\in T_{o_{n}}H^{n}$. Since $f_{2}^{-1}$ is injective and $\mathbf{v}\neq\mathbf{0}$, $\mathbf{u}\neq\mathbf{0}$. Let $x=f_{1}^{-1}(\mathbf{u})\in H^{n}$. Since $f_{1}^{-1}$ is injective (it is $\exp_{o_{n}}^{c}$) and $\mathbf{u}\neq\mathbf{0}$, $x\neq o_{n}$. Consider $F(x)$ and $F(o_{n})$. $F(x)=f_{5}(f_{4}(f_{3}(f_{2}(f_{1}(x)))))=f_{5}(f_{4}(f_{3}(\mathbf{u})))=f_{5}(f_{4}(\mathbf{W}\mathbf{v}))=f_{5}(f_{4}(\mathbf{0}))$. $f_{4}(\mathbf{0})=(0,\mathbf{0})\in T_{o_{d}}H^{d}$. $F(x)=\exp_{o_{d}}^{c}(0,\mathbf{0})$. The exponential map of the zero vector in the tangent space at $p$ is always $p$. So, $\exp_{o_{d}}^{c}(0,\mathbf{0})=o_{d}$. Now consider $F(o_{n})$: $F(o_{n})=\exp_{o_{d}}^{c}\left(0,\mathbf{W}\left(\log_{o_{n}}^{c}o_{n}\right)_{[1:n]}\right)$. $\log_{o_{n}}^{c}o_{n}$ is the zero vector in $T_{o_{n}}H^{n}$, which is $(0,\mathbf{0})\in\mathbb{R}^{n+1}$. $\left(\log_{o_{n}}^{c}o_{n}\right)_{[1:n]}=\mathbf{0}\in\mathbb{R}^{n}$. $F(o_{n})=\exp_{o_{d}}^{c}(0,\mathbf{W}\mathbf{0})=\exp_{o_{d}}^{c}(0,\mathbf{0})=o_{d}$. So $F(x)=o_{d}$ and $F(o_{n})=o_{d}$. Since $x\neq o_{n}$, we have found two distinct points $x,o_{n}\in H^{n}$ that map to the same point $o_{d}\in H^{d}$. Thus, $F$ is not injective if $\mathbf{W}$ does not have full column rank.